Math Homework

Nipride 3 mcg/kg/min has been ordered for a patient who weighs 192 pounds. The solution is labeled 50 mg in 250 mL of D5W.

a. What is your dosage per minute? ________________________________________

b. What is your mL per hour flow rate? ____________
Answer #1

192 pounds = 87 kgs

Nipride 3 mcg/ kg / min = 3 mcg \times 87 = 261 mcg

a. The dosage per minute is 261 mcg or 0.26 mg

b. For 1 hour the dosage is 261 \times 60 = 15660 mcg or 15.66 mg

Required medication = Required dose \times available volume \div available dose

                                  = 15.66 \times 250 \div 50

                                  = 78.3 or 78

The flow rate is 78.3 or 78 ml per hour